Please explain this GDP problem? Why do you/don't you include the things you include? Answer is C ($575).

30. Table-8: Calculate the GDP of the country described in the table. Note that there are no income or expenditures other than those described in the table A $825 B $650 Interest Income Wages Business Profits Rental Income Household Investment Household Consumption Government Investment Other Government Expenditures and Transfers Imports Exports $100 $250 $125 $100 $175 $300 $125 $150 $75 $50 Table-8 C $575 D $1,150
Answer #1

GDP=C+I+G+NX

We include all the consumptions, investments and net exports. But, we should not include the transfers.

GDP = 175 + 300 + 125 + (50-75) = 575

(Explanation why transfers are not included - transfers are done as subsidy or scholarship or similar schemes. Unlike payment, these are not in exchange of any good or service. Hence a transfer do not increase the real income or output in the economy. It merely transfers income from one hand to another)